Project

General

Profile

Actions

2016-01-08-Physics

Physics coordinator report, 2016-01-08

  • Optics review
    • Paperwork has been put together, a bit too long, Ken is doing some cutting
    • Talks due on Monday for practice session
  • SSD w/o M1
    • Two new iterations from Ao on settings (17 December and 31 December), following some back and forth about current limits and other details
    • we have settled on a set for the optics review;
    • Bz looks rather low in d/s spectrometer (~ 2 T)
    • analysis of tracker reconstruction (Rogers, Dobbs) in progress
  • Detector alignment
    • Francois, Chris Hunt, both leant on to produce notes
    • John Nugent has successful MC alignment
  • Magnet alignment
    • Sophie is concentrating on transfer matrix-based FC alignment; validating routines using straight track data (where analytical transfer matrix is obvious)
  • PID
    • TOF vs tracker PID first pass has been done; 2 % pion impurity under the muon peak (4 T run) looks quite comparable with pion contamination paper
  • Xenon analysis
    • Data taken, first pass analysis performed, comparison with no absorber data ongoing
    • Rumours/discovery of Helium in "straight tracks" absorber; who is responsible for material budget?
  • EMR paper accepted by J. Inst.
  • Pion contamination paper - J. Inst. struggling to find referees
  • Demo paper - no progress
  • Step IV paper - see VC report
    • Next step is to make latex outline of the MICE note and organise deadlines
    • Expect a couple of iterations on tracker reconstruction at least

Updated by Rogers, Chris over 7 years ago ยท 1 revisions